OpenStudy (anonymous):

position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

geerky42 (geerky42):

\[\large v(t) = \dfrac{\mathrm d }{\mathrm dt}~s(t)\]Right?

geerky42 (geerky42):

You just need to evaluate \(v(2)\)

OpenStudy (anonymous):

If I evaluate the formula at t=-2 I just get -14

OpenStudy (anonymous):

*2

geerky42 (geerky42):

How did you get -14? Can you find derivative of -2 - 6t ?

OpenStudy (anonymous):

I don't think I know how to find derivatives properly.

OpenStudy (anonymous):

When it says derivative, I think derived from t equaling 2 and the formula -2 -6t. -2-6*(2)=-14

OpenStudy (anonymous):

I don't think that's how it works though.

geerky42 (geerky42):

You can just take derivative of each term individually. For each term, use power rule: \(\dfrac{\mathrm d }{\mathrm dx}x^n = nx^{n-1}\) For constant, you have \(\dfrac{\mathrm d }{\mathrm dx}c = 0\)

geerky42 (geerky42):

So here, you have \(\dfrac{\mathrm d }{\mathrm dt}(-2-6t) ~~~=~~~ \dfrac{\mathrm d }{\mathrm dt}(-2) + \dfrac{\mathrm d }{\mathrm dt}(-6t) = ~?\)

OpenStudy (anonymous):

(d/-2d)-(d/-2d)=0

OpenStudy (anonymous):

?

geerky42 (geerky42):

Seem you are not really familiar with basic derivative rules. Can you tell me what you learned recently (what teacher tried to teach you, etc)?

OpenStudy (anonymous):

I don't have a teacher

OpenStudy (anonymous):

Can you tell me what you got, then we can work it out.

geerky42 (geerky42):

So you are self-taught?

OpenStudy (anonymous):

Probably online school :) ^^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If someone could offer me the answer to this question, I would be glad to discuss how it was found